the width of a rectangle is 6 less than twice the length, if the perimeter is 18 meters.find the length
Answer by nerdybill(7384)

the width of a rectangle is 6 less than twice the length, if the perimeter is 18 meters.find the length
.
Let x = length
then
2x-6 = width
.
2(x + 2x-6) = 18
(x + 2x-6) = 9
3x-6 = 9
3x = 15
x = 5 meters (length of rectangle)